OpenStudy (anonymous):

Find the critical value z a/2 that corresponds to a 94% confidence level

OpenStudy (anonymous):

You have an alpha level of 0.06 a/2 = 0.06/2 = 0.03 0.05 - 0.03 = 0.4700 now look up the z value for p=0.4700 in the z-table and thats your answer.